# Return of the Lepton Number: Sterile Neutrino Dark Matter Production and the Revival of the Shi-Fuller Mechanism

### v1

@@ III Results
Our primary results are shown in Figure 6, which was calculated by using the GZA code as described in the previous section over the full parameter space shown. The colored contour plot shows the equivalent thermal WDM particle mass for that position in parameter space. The white dot-dashed line at $m_{\mathrm{thermal}}=4.1\,\mathrm{keV}$ is the best-fit value for the evidence for WDM from the Lyman-$\alpha$ forest [20], and the full parameter space is consistent with the $1\sigma$ inferred thermal WDM particle mass from that work. In Fig. 6, the red dashed line represents the current strongest bound on thermal WDM particle mass $m_{\mathrm{th}}>8.5\,\mathrm{keV}$ (95% CL) from a combination of galaxy counts [18] and strong lensing (Zelko+) [19]. The gray exclusion region is formed from the strongest X-ray bound at each mass, resulting from NuSTAR and SPI/INTEGRAL constraints [39, 40, 41, 42, 43].
The contours for constant $L$ shift down with increasing $m_{s}$ because a smaller number density is needed to match the observed dark matter density, and less $L$-driven resonant conversion is needed. We find that the available window for SF to account for the entirety of the dark matter is excluded at much greater than $95\%\,\mathrm{CL}$ for $L<0.1$; however, extending to higher $L$ demonstrates the potential for a much larger parameter space available based on the current large values possible for the lepton number in the early Universe. The shapes for the contours of constant $m_{\mathrm{th}}$ have a tilt up and to the left due to the fact, as shown in Fig. 1, that higher $L$ in this region of the parameter space leads to population of higher $q$ momenta modes, and effectively warmer dark matter. Note that at higher mixing angles, in the regions shown that are largely ruled out by X-ray observations, the contours reverse as the SF mechanism produces lower average $q$ distributions, as shown in Ref. [12].
@@
$m_{\mathrm{th}}$ $=A\,m_{s}^{a}\,\left[\sin^{2}(2\theta)\right]^{b},\quad\begin{cases}A=1.329\\ a=1.044\\ b=0.055\,,\end{cases}$ (9)
$L$ $=C\,m_{s}^{\alpha}\,\left[\sin^{2}(2\theta)\right]^{\beta},\quad\begin{cases}C=1.865\times 10^{-12}\\ \alpha=-1.813\\ \beta=-1.042\,,\end{cases}$ (10)
where particle masses are in units of keV. Importantly, $m_{\mathrm{th}}$ in these relations is the more accurate thermal WDM particle mass calculated in Ref. [21].
These expressions capture the dominant scaling behavior in the high-lepton asymmetry regime and are validated to remain within percent-level error bounds across the range $15~{}\mathrm{keV}\leq m_{s}\leq 60~{}\mathrm{keV}$ and $10^{-15}\leq\sin^{2}(2\theta)\leq 3\times 10^{-13}$. Specifically, the average error for the thermal mass fit was 1.3% while the average error for the lepton number fit was 8.3%.
Interestingly, if one adopts the Zelko+ bounds, the remaining unconstrained parameter space requires an $L\gtrsim 0.5$. Note that the full parameter space in Fig. 6 is preferred by Lyman-$\alpha$ observations and within the derived $1\sigma$ thermal WDM particle mass region given by $m_{\mathrm{th}}=4.1^{+33.3}_{-1.5}\,\mathrm{keV}$ [20].
@@ IV Discussion and Conclusions
In this paper, we examine the Shi-Fuller sterile neutrino dark matter production mechanism given the newly uncovered possibility of large lepton numbers in the early Universe, during the dark matter production era. In this context, we study the combined constraints from structure formation and X-ray astronomy. Interestingly, there are both structure formation constraints on WDM from lensing and galaxy counts, as well as a current preference from recent interpretations of Lyman-$\alpha$ forest measurements. Specifically, the thermal WDM particle mass preferred by an interpretation of Lyman-$\alpha$ forest observations is $m_{\mathrm{th}}=4.1^{+33.3}_{-1.5}\,\mathrm{keV}$ [20], while constraints on WDM from galaxy counts and lensing are at the level of $m_{\mathrm{th}}>8.4\,\mathrm{keV}$ (95% CL) [18] and $m_{\mathrm{th}}>8.5\,\mathrm{keV}$ (95% CL) [19]. [^4] As discussed in Sec. I, we use the updated, more accurate thermal WDM transfer functions for these signals and constraints. These structure formation constraints and preferences lie in the parameter space for SF sterile neutrino dark matter production that are not constrained by X-ray astronomy. We find that the available parameter space for the SF mechanism is no longer closed, but opens up for $L\gtrsim 0.5$, $m_{s}\gtrsim 35\,\mathrm{keV}$, and $\sin^{2}(2\theta)\lesssim 10^{-14}$ when adopting both structure formation and X-ray constraints.
@@
There are several other observational and experimental probes of the SF mechanism for the production of sterile neutrino dark matter. One is a potential detection of $N_{\mathrm{eff}}>3.044$ with high significance from upcoming CMB experiments [83, 82]. This could be a signature of high-$L$ in the very early Universe [54]. $N_{\mathrm{eff}}\approx 3.5$ is already favored by the Hubble tension [86, 87]. As discussed in Section II.2, the primordial abundances are also very sensitive to non-zero lepton number (chemical potential), particularly in the electron-flavor neutrinos. Of interest specifically is the the recent EMPRESS measurement of $Y_{p}$ that may indicate nonzero $L_{e}$ at the BBN epoch [88, 89, 90, 91], which would expand the possibilities for larger $L_{\alpha}$ in all flavors at early times [54]. CMB observations are also becoming highly sensitive to $Y_{p}$ and $N_{\mathrm{eff}}$ independently, allowing for tests of primordial nucleosynthesis [92, 93].
@@
Since the parameter space remaining is at lepton numbers of order unity, we have identified approximations that must be examined in this high-$L$ regime of early Universe physics. In the fortunate event that a future X-ray observatory detects a signal in this small mixing angle region, our work shows what parameter spaces are of interest for structure formation and potential nuclear decay probes [94, 95, 96, 97], as well as the necessity of carefully examining the full quantum kinetic behavior of SF sterile neutrino dark matter production at high-$L$.
@@
In addition to the resonance of the SF mechanism, enhancing the active neutrinos’ scattering rate through nonstandard interactions (so-called NSI) can allow access to $\sin^{2}(2\theta)\gtrsim 10^{-14}$ for sterile neutrino dark matter production via oscillation-based production [98, 99, 100, 101]; however, this NSI mechanism is beyond the scope of this present paper, and will be examined in more detail in a forthcoming paper [102]. In addition to enhancing methods of production involving neutrino oscillations, sterile neutrino dark matter can also be produced by mechanisms that are independent of mixing, such as those from parent particle decay to sterile neutrinos, including the “keV miracle” from Higgs boson decay [103], a GUT-scale mechanism [104], vector boson decay [105], from thermal production and subsequent dilution by new particles [106], from a parent sterile neutrino [107], or via Yukawa couplings [108].
@@ Acknowledgements.
CMV, KNA and HGE acknowledge useful conversations with Francis-Yan Cyr-Racine, George Fuller, Jay Krishnan, Harri Parkkinen, Michael Ryan, and M. Cristina Volpe. JF thanks Cyril Pitrou for many useful discussions. KNA is partially supported by U.S. National Science Foundation (NSF) Theoretical Physics Program Grant No. PHY-2210283. KNA acknowledges support of the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université), and LabEx CARMIN (ANR-10-LABX-59-01) in hosting the “Dark Matter and Neutrinos” program, where many of these discussions took place. JF is supported by the Network for Neutrinos, Nuclear Astrophysics and Symmetries (N3AS), through the National Science Foundation Physics Frontier Center Award No. PHY-2020275.

### v3

@@ III Results
Our primary results are shown in Fig. 7, which was calculated by using the GZA code as described in the previous section over the full parameter space shown. The colored contour plot shows the equivalent thermal WDM particle mass for that position in parameter space. The white dot-dashed line at $m_{\mathrm{thermal}}=4.1\,\mathrm{keV}$ is the best-fit value for the evidence for WDM from the Lyman-$\alpha$ forest [Villasenor:2022aiy], and the full parameter space is consistent with the $1\sigma$ inferred thermal WDM particle mass from that work. In Fig. 7, the red dashed line represents the current strongest bound on thermal WDM particle mass $m_{\mathrm{th}}>8.5\,\mathrm{keV}$ (95% CL) from a combination of galaxy counts [Nadler:2021dft] and strong lensing (Zelko+) [Zelko:2022tgf]. The gray exclusion region is formed from the strongest x-ray bound at each mass, resulting from NuSTAR and SPI/INTEGRAL constraints [Perez:2016tcq, Roach:2019ctw, Roach:2022lgo, Krivonos:2024yvm, Fischer:2022pse].
The contours for constant $L$ shift down with increasing $m_{s}$ because a smaller number density is needed to match the observed dark matter density, and less $L$-driven resonant conversion is needed. We find that the available window for SF to account for the entirety of the dark matter is excluded at much greater than $95\%\,\mathrm{CL}$ for $L<0.1$; however, extending to higher $L$ demonstrates the potential for a much larger parameter space available based on the current large values possible for the lepton number in the early Universe. The shapes for the contours of constant $m_{\mathrm{th}}$ have a tilt up and to the left due to the fact, as shown in Fig. 1, that higher $L$ in this region of the parameter space leads to a population of higher $q$ momenta modes, and effectively warmer dark matter. Note that at higher mixing angles, in the regions shown that are largely ruled out by x-ray observations, the contours reverse as the SF mechanism produces lower average $q$ distributions, as shown in Ref. [Cherry:2017dwu].
@@
[…]
where particle masses are in units of keV. Importantly, $m_{\mathrm{th}}$ in these relations is the more accurate thermal WDM particle mass calculated in Ref. [Vogel:2022odl].
These expressions capture the dominant scaling behavior in the high-lepton asymmetry regime and are validated to remain within percentage-level error bounds across the range $15~\mathrm{keV}\leq m_{s}\leq 60~\mathrm{keV}$ and $10^{-15}\leq\sin^{2}(2\theta)\leq 3\times 10^{-13}$. Specifically, the average error for the thermal mass fit was 1.3% while the average error for the lepton number fit was 8.3%.
Interestingly, if one adopts the Zelko+ bounds, the remaining unconstrained parameter space requires an $L\gtrsim 0.5$. Note that the full parameter space in Fig. 7 is preferred by Lyman-$\alpha$ observations and within the derived $1\sigma$ thermal WDM particle mass region given by $m_{\mathrm{th}}=4.1^{+33.3}_{-1.5}\,\mathrm{keV}$ [Villasenor:2022aiy].
@@ IV Discussion and Conclusions
In this paper, we examine the Shi-Fuller sterile neutrino dark matter production mechanism given the newly uncovered possibility of large lepton numbers in the early Universe, during the dark matter production era. In this context, we study the combined constraints from structure formation and x-ray astronomy. Interestingly, there are both structure formation constraints on WDM from lensing and galaxy counts, as well as a current preference from recent interpretations of Lyman-$\alpha$ forest measurements. Specifically, the thermal WDM particle mass preferred by an interpretation of Lyman-$\alpha$ forest observations is $m_{\mathrm{th}}=4.1^{+33.3}_{-1.5}\,\mathrm{keV}$ [Villasenor:2022aiy], while constraints on WDM from galaxy counts and lensing are at the level of $m_{\mathrm{th}}>8.4\,\mathrm{keV}$ (95% CL) [Nadler:2021dft] and $m_{\mathrm{th}}>8.5\,\mathrm{keV}$ (95% CL) [Zelko:2022tgf]. [^4] As discussed in Sec. I, we use the updated, more accurate thermal WDM transfer functions for these signals and constraints. These structure formation constraints and preferences lie in the parameter space for SF sterile neutrino dark matter production that are not constrained by x-ray astronomy. We find that the available parameter space for the SF mechanism is no longer closed, but opens up for $L\gtrsim 0.5$, $m_{s}\gtrsim 35\,\mathrm{keV}$, and $\sin^{2}(2\theta)\lesssim 10^{-14}$ when adopting both structure formation and x-ray constraints.
@@
There are several other observational and experimental probes of the SF mechanism for the production of sterile neutrino dark matter. One is a potential detection of $N_{\mathrm{eff}}>3.044$ with high significance from upcoming CMB experiments [CMB-S4:2016ple, SimonsObservatory:2025wwn]. This could be a signature of high-$L$ in the very early Universe [Froustey:2024mgf]. $N_{\mathrm{eff}}\approx 3.5$ is already favored by the Hubble tension [Escudero:2022rbq, Escudero:2024uea]. As discussed in Sec. II.2, the primordial abundances are also very sensitive to nonzero lepton number (chemical potential), particularly in the electron-flavor neutrinos. Of interest specifically is the recent EMPRESS measurement of $Y_{p}$ that may indicate nonzero $L_{e}$ at the BBN epoch [Matsumoto:2022tlr, Burns:2022hkq, Escudero:2022okz, Yanagisawa:2025mgx], which would expand the possibilities for larger $L_{\alpha}$ in all flavors at early times [Froustey:2024mgf]. CMB observations are also becoming highly sensitive to $Y_{p}$ and $N_{\mathrm{eff}}$ independently, allowing for tests of primordial nucleosynthesis [ACT:2025tim, SPT-3G:2025bzu].
@@
Since the parameter space remaining is at lepton numbers of order unity, we have identified approximations that must be examined in this high-$L$ regime of early Universe physics. In the fortunate event that a future x-ray observatory detects a signal in this small mixing angle region, our work shows what parameter spaces are of interest for structure formation and potential nuclear decay probes [Friedrich:2020nze, Martoff:2021vxp, KATRIN:2022spi, Lee:2025txi], as well as the necessity of carefully examining the full quantum kinetic behavior of SF sterile neutrino dark matter production at high $L$.
@@
In addition to the resonance of the SF mechanism, enhancing the active neutrinos’ scattering rate through nonstandard interactions (so-called NSI) can allow access to $\sin^{2}(2\theta)\gtrsim 10^{-14}$ for sterile neutrino dark matter production via oscillation-based production [DeGouvea:2019wpf, Kelly:2020pcy, Berryman:2022hds, An:2023mkf]; however, this NSI mechanism is beyond the scope of this present paper, and was recently examined in more detail in Ref. [Vogel:2025gan]. In addition to enhancing methods of production involving neutrino oscillations, sterile neutrino dark matter can also be produced by mechanisms that are independent of mixing, such as those from parent particle decay to sterile neutrinos, including the “keV miracle” from Higgs boson decay [Kusenko:2006rh], a Grand-Unified-Theory-scale mechanism [Kusenko:2010ik], vector boson decay [Shuve:2014doa], from thermal production and subsequent dilution by new particles [Patwardhan:2015kga], from a parent sterile neutrino [Fuller:2024noz], or from new active-sterile interactions [Dev:2025sah], including via varying Yukawa couplings [Jaramillo:2022mos].
@@
Note added. — Recently, a related paper appeared on arXiv [Akita:2025txo]. In that paper, the authors find similar expansion of the parameter space using a new production code which incorporates nonaveraged neutrino oscillations and additional effects from large chemical potentials.
@@ Acknowledgements.
C.M.V., K.N.A. and H.G.E. acknowledge useful conversations with Francis-Yan Cyr-Racine, George Fuller, Jay Krishnan, Harri Parkkinen, Michael Ryan, and M. Cristina Volpe. J.F. thanks Cyril Pitrou for many useful discussions. K.N.A. is partially supported by U.S. National Science Foundation (NSF) Theoretical Physics Program Grant No. PHY-2210283. K.N.A. acknowledges support of the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université), and LabEx CARMIN (ANR-10-LABX-59-01) in hosting the “Dark Matter and Neutrinos” program, where many of these discussions took place. J.F. is supported by the Network for Neutrinos, Nuclear Astrophysics and Symmetries (N3AS), through the National Science Foundation Physics Frontier Center Award No. PHY-2020275.